OpenStudy (babyslapmafro):

[cos^2(x)]^(3/2)=cos^3(x) correct? (click to see a more clear example in equation-format)

OpenStudy (babyslapmafro):

\[(\cos ^{2}x) ^{\frac{ 3 }{ 2 }}=\cos ^{3}x\]

OpenStudy (babyslapmafro):

correct?